The Ramirez Company's last dividend was $1.75. Its dividend growth rate is expected to be constant at 25% for 1 year, after which dividends are expected to grow at a rate of 6% forever. Its required return (rs) is 8%. What is the best estimate of the current stock price?

-$120.11

-$98.44

-$115.29

-$103.22

-$109.38
Answer #1

D1=(1.75*1.25)=2.1875

Value after year 1=(D1*Growth rate)/(Required return-Growth rate)

=(2.1875*1.06)/(0.08-0.06)

=115.9375

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.1875/1.08+115.9375/1.08

=$109.38(Approx).